An open box is made from a square piece of material 24 inches on a side by cutting equal squares from the corners and turning up the sides. Write the Volume V of the box as a function of x. Recall that Volume is the product of length, width, and height.

Thank you!
Answer #1
Or V=1/3(24-2x)^2*x

I'm confused as to how you came up with the base as 24-2x.
Thanks for your help!
answered by: neeta
Answer #2
V(x)= x(24-x)^2 ?
answered by: jjr
Answer #3
draw a square with sides 24 in
draw small square cut-outs at each of the corners, each of those sides is x inches long.
Now didn't you take away x inches from each end??
so 24 - x -x or 24-2x
answered by: Candyce
Answer #4
let each side of the cut-out square be x in
then isn't the base 24-2x by 24-2x ?
and isn't the height of the box x ?

take it from there.
answered by: lena Hose
Answer #5
Okay, I just forgot the 2. Thank you so much for your help! Have a nice day!
answered by: Princess
Answer #6
How did you come up with that equation ??

Your equation of V(x)= x(24-x)^2
in your first response should have been

V(x)= x(24-2x)^2 ?
answered by: UCI Student